OpenStudy (anonymous):

Please help me with trig: sec((5pi)/3)=?

OpenStudy (anonymous):

\[\huge \sec\left(\frac{5\pi}{3}\right)=\frac{1}{\cos\left(\frac{5\pi}{3}\right)}\]

OpenStudy (anonymous):

and cos 5pi/3 is sqrt2/2, right?

OpenStudy (anonymous):

No, it isn't :)

OpenStudy (anonymous):

that's what i saw on the trig circle. what is it?

OpenStudy (anonymous):

Actually, it's 1/2 :)

OpenStudy (anonymous):

that sounds better, but how do you figure?

OpenStudy (anonymous):

bc when i use my calculator i get 2 as the answer

OpenStudy (anonymous):

That's because it's COSINE that's 1/2 But you're asked for SECANT, which is 1/Cosine ^.^

OpenStudy (anonymous):

yeah, i got that part.. i'm confused as to how cosine5pi/3 equals 1/2 and not sqrt2/2

OpenStudy (anonymous):

nevermind! i was reading the spot next to it

OpenStudy (anonymous):

thank you very much:)